# Theory of Mind Using Active Inference: A Framework for Multi-Agent Cooperation

### v1

@@ Abstract
We present a novel approach to multi-agent cooperation by implementing theory of mind (ToM) within active inference. ToM – the ability to understand that others can have differing knowledge and goals – enables agents to reason about others’ beliefs while planning their own actions. Unlike previous active inference approaches to multi-agent cooperation, our method neither relies on task-specific shared generative models nor requires explicit communication, while being generalisable. In our framework, the ToM-equipped agent maintains distinct representations of its own and others’ beliefs and goals. We extend the sophisticated inference tree-based planning algorithm to systematically explore joint policy spaces through recursive reasoning. Our approach is evaluated through collision avoidance and foraging task simulations. Results demonstrate that ToM-equipped agents cooperate better compared to non-ToM counterparts by being able to avoid collisions and reduce redundant efforts. Crucially, ToM agents accomplish this by inferring others’ beliefs solely from observable behaviour. This work advances practical applications in artificial intelligence while providing computational insights into ToM.
@@ 1 Introduction
1.1 Multi-Agent Cooperation using Active Inference
Maisto and colleagues [9] introduced “interactive inference”, wherein agents maintain probabilistic beliefs about shared goals (such as both agents pressing the same or a different button) and update these beliefs through observations of others’ locations and actions. Their agents selected epistemic policies designed to reduce uncertainty about the joint goals. However, this approach assumes agents share identical goals, which is not always the case in multi-agent cooperation tasks. Moreover, their model relies on a carefully tailored generative model for the focal agent (i.e., the agent conducting ToM) that incorporates the other agent’s location as an observation that reveals information about the shared goal. These assumptions limit generalisability to scenarios where actions do not inherently signal goals or where agents have complementary rather than identical and shared objectives.
Matsumura and colleagues [10] addressed collision avoidance (agents passing by each other without colliding) using simulation theory, where agents use their own internal models to imagine others’ situations. While this enables basic perspective-taking, their implementation is domain-specific to navigation tasks that use the social force model, with parameters for forward movement and mutual repulsion. The approach lacks the recursive reasoning capabilities that relies on maintenance of separate belief representations for different agents, which is necessary for more complex coordination scenarios.
@@ 1.1 Multi-Agent Cooperation using Active Inference
Overall, these approaches predominantly assume that all agents operate under the same generative model, with identical beliefs about transition dynamics, observation likelihoods and goal structures. Such aligned models fail to capture the reality of agents with different experiences, capabilities, and objectives. Furthermore, these approaches typically involve only single-level reasoning (“what will the other agent do?”) rather than the recursive beliefs that characterise ToM (“what do I think the other agent thinks about the situation?”). Many implementations are also tailored to specific tasks (e.g., navigation, mutual button-pressing) and do not provide general principles for multi-agent cooperation across different domains.
@@
    Our agents can reason about how others’ actions affect the world states via likelihood message passing between their own beliefs and their beliefs of the other agent’s beliefs, maintaining perspective separation while allowing for information integration.
@@
2 Theory of Mind in Active Inference
@@ 2.2 ToM Agent’s Belief Structure of Multiple Agents
In our ToM framework, the focal agent maintains separate beliefs for itself and each other agent in the environment. In the case of a two-agent scenario, this yields $s=\{\underbrace{s^{f,\text{self}},s^{f,\text{world}}}_{\text{focal ($s^{f}$)}},\underbrace{s^{o,\text{self}},s^{o,\text{world}}}_{\text{other ($s^{o}$)}}\}$, where:
@@
This structure enables the focal agent to maintain distinct perspectives while representing how others’ beliefs might differ from its own. To elaborate, these belief components can be flexibly combined to model different reasoning scenarios. For instance, the focal agent can pair $s^{o,\text{self}}$ with $s^{f,\text{world}}$ to predict what the other agent would observe from its current location based on the focal agent’s own understanding of the world states. This cross-perspective reasoning allows the focal agent to distinguish between what it believes the other agent knows versus what the other agent would perceive given the focal agent’s view of reality. Crucially, the focal agent’s beliefs about the world states, and its beliefs about the other agent’s beliefs about these same states, can diverge. This allows the agent to represent scenarios where it possesses knowledge that the other agent lacks, or vice versa.
@@ 2.3 Recursive Planning with Theory of Mind
Step 1: Other Agent Policy Expansion. As aforementioned in Section 2.2, we begin with the focal agent’s beliefs, which comprises separate beliefs for itself and the other agent in the environment ($s=\{s^{f},s^{o}\}$). The focal agent first considers which policies the other agent is likely to select. This is visualised in Step 1 of Figure 1, where each policy node represents a specific action that the other agent might execute ($a_{0}$; purple square). The potential actions are evaluated based on the focal agent’s beliefs about the other agent’s beliefs ($s^{o}$; purple circle above the purple square). Essentially, the focal agent asks “Given what I believe about the other agent’s beliefs and goals, what would the other agent choose to do?”. The focal agent then computes how the other agent would update its beliefs if it were to execute that action ($s^{o}$; the purple circle below the purple square).
@@
Step 2: Focal Agent Policy Expansion. For each considered action of the other agent, the focal agent evaluates its own policy options. Crucially, before doing so, the focal agent updates its world beliefs, based on the anticipated consequences of the other agent’s actions, using likelihood message passing [2]. Here, the focal agent uses its computation of how the other agent’s beliefs about the world ($s^{o,\text{world}}$; purple circle above the red circle) would change if the other agent were to execute an action. A likelihood message is then created, capturing the information gained from the other agent’s anticipated action in the form of the difference between the other agent’s updated beliefs and its prior beliefs. This mechanism allows the focal agent to incorporate information into its own beliefs about how the world states ($s^{f,\text{world}}$; red circle above the red squares) would change resulting from the other agent’s actions. Using these updated beliefs, the focal agent then evaluates its own policy options ($a_{0}$; red square) through standard EFE calculations. This creates branches in the tree structure for each possible joint policy combination between the focal and other agent. The focal agent then computes how its beliefs would be updated if it were to execute an action given the other agent’s action ($s^{f}$; red circle below the red square).
Step 3: Other Agent Observation Expansion. The focal agent considers what observations the other agent ($o_{0}$; purple diamonds) is likely to receive given the joint policy. These observation probabilities are computed using the focal agent’s beliefs about the other agent’s self states ($s^{o,\text{self}}$; purple circle from an earlier expansion) and the focal agent’s own updated beliefs about the world state ($s^{f,\text{world}}$; red circle before the purple diamonds). The focal agent then updates its representation of the other agent’s posterior beliefs ($s^{o}$; purple circle after the purple diamond).
@@
Step 4: Focal Agent Observation Expansion. Then, the focal agent considers likely observations and resulting posterior beliefs given the joint policy and anticipated world state changes. At this stage, the focal agent updates its beliefs about both itself and the other agent.
@@
This process is illustrated in Step 4 of Figure 1, where the focal agent’s observation nodes ($o_{0}$; red diamonds) represent the various observations the focal agent expects to encounter given the focal agent’s beliefs ($s^{f}$; red circle before red diamonds) considering the execution of both agents’ actions. This results in the computation of the focal agent’s posterior beliefs ($s^{f}$; red circle after red diamonds), which reflects the updated beliefs that incorporate both its own and the other agent’s perspective.
@@ 3.1 Collision Avoidance Task
Preferences encoded goal-seeking behaviour: high positive utility for reaching the target location and severe penalty for the null observation. Critically, no explicit collision avoidance preferences were included – coordination had to emerge through ToM reasoning rather than hard-coded behaviours. Planning horizons were set to 4 time steps for both ToM and non-ToM agents, sufficient to reach goals via alternative paths while requiring that much forward planning to identify coordination opportunities.
@@
In the ToM condition, the red (ToM) agent reasoned that the other agent will most likely move to the central location to take the shortest path to its goal, so it does not select moving to that location to avoid collision even though it was the most optimal policy for itself. Instead, the agent went around the centre, which was the next best alternative route towards its goal. The ToM agent selected a longer but collision-free route.
@@ 3.2 Apple Foraging Task
The environment was partially observable such that agents could only assess apple availability at their current location, creating uncertainty about the broader resource distribution. Apples spawned probabilistically in orchard locations (25% per time step) and remained in that location until consumption. Action repertoires included movements (up, down, left, right actions), eating, and no operation. Preferences simply favoured reward acquisition without explicit cooperation incentives. Planning horizons were set to 2 time steps, sufficient for reaching adjacent locations and consuming apples.
@@
For an illustration of the results from this task, see Figure 2(b). For the detailed planning tree of the ToM vs non-ToM agents at time step 0, see Appendix 0.B.
@@ Appendix 0.A Expected Free Energy under Sophisticated Inference with Theory of Mind
Appendix 0.B Planning Trees for Apple Foraging Task
Figure 3 shows some detailed planning trees for both a non-ToM and ToM agent, highlighting the differences on a concrete example. We start from the orchard environment state in (a), where both agents navigate to locate and consume apples, with initial shared knowledge of an apple at location 9. We visualise the planning tree for the red agent, for both a non-ToM (c) and ToM (d) variant.
In the non-ToM case, the focal agent (red) evaluates only its own policies over a 2-step horizon, selecting to go to location 9 (P$=$1.0) based on expected utility (G$=$10.00). However, when pursuing this policy it will end up with no apple, as the purple agent will get there first.
@@ Appendix 0.B Planning Trees for Apple Foraging Task
Figure 3: (a) Start situation with the red agent as the focal agent. (b) Reference grid layout for 3×3 simulation environment. Numbers 1-9 indicate cell indices used throughout the experiments to specify agent locations and movements. (c) Non-ToM planning tree: The red agent evaluates only its own policies over a 2-step horizon, selecting to go to location 9 (P$=$1.0) based on expected utility (G$=$10.00). (d) ToM planning tree: The red (ToM) agent recursively models the purple (non-ToM) agent’s policy space and beliefs, resulting in selection of the third policy with full certainty (q_pi$=$1.0) which is to go left. This recursive reasoning leads to exploration of location 7, as the agent predicts the purple agent will pursue location 9, enabling coordinated resource allocation.

### v2

@@ Abstract
Theory of Mind (ToM) – the ability to understand that others can have differing knowledge and goals – enables agents to reason about others’ beliefs while planning their own actions. We present a novel approach to multi-agent cooperation by implementing ToM within active inference. Unlike previous active inference approaches to multi-agent cooperation, our method neither relies on task-specific shared generative models nor requires explicit communication. In our framework, ToM-equipped agents maintain distinct representations of their own and others’ beliefs and goals. ToM agents then use an extended and adapted version of the sophisticated inference tree-based planning algorithm to systematically explore joint policy spaces through recursive reasoning. We evaluate our approach through collision avoidance and foraging simulations. Results suggest that ToM agents cooperate better compared to non-ToM counterparts by being able to avoid collisions and reduce redundant efforts. Crucially, ToM agents accomplish this by inferring others’ beliefs solely from observable behaviour and considering them when planning their own actions. Our approach shows potential for generalisable and scalable multi-agent systems while providing computational insights into ToM mechanisms.
@@ 1 Introduction
1.1 Existing Active Inference Approaches to Multi-Agent Cooperation
Maisto and colleagues [9] introduced “interactive inference”, wherein agents maintain probabilistic beliefs about shared goals (such as both agents pressing the same or different buttons) and update these beliefs through observations of others’ locations and actions. Their agents selected epistemic policies designed to reduce uncertainty about the joint goals. However, this approach assumes agents share identical goals, which is not always the case in multi-agent cooperation tasks. Moreover, their model relies on a carefully tailored generative model for the focal agent (i.e., the agent conducting ToM) that incorporates the other agent’s location as an observation which itself encodes information about the shared goal. These assumptions limit generalisability to scenarios where actions do not inherently signal goals or where agents have complementary rather than identical and shared objectives.
Matsumura and colleagues [10] addressed collision avoidance (agents passing by each other without colliding) using simulation theory, where agents use their own internal models to imagine others’ situations. While this enables basic perspective-taking, their implementation is domain-specific to navigation tasks that use the social force model as it includes parameters for forward movement and mutual repulsion. The approach lacks the recursive reasoning capabilities that relies on maintenance of separate belief representations for different agents, which is necessary for more complex coordination scenarios.
@@ 1.1 Existing Active Inference Approaches to Multi-Agent Cooperation
Overall, these approaches predominantly assume that all agents operate under the same generative model, with identical beliefs about transition dynamics, observation likelihoods and goal structures. Such aligned models fail to capture the reality of agents with different experiences, capabilities, and objectives. Furthermore, these approaches typically involve only single-level reasoning (“what will the other agent do?”) rather than the recursive beliefs that characterise ToM (“what do I think the other agent thinks about the situation?”). Many implementations are also tailored to specific tasks (e.g., navigation or mutual button-pressing) and do not provide general principles for multi-agent cooperation across different tasks.
@@
    Our agents can reason about how others’ actions affect the world states via message passing between their own beliefs and their beliefs of the other agent’s beliefs, maintaining perspective separation while allowing for information integration.
@@
2 Our Approach: Theory of Mind in Active Inference
@@ 2.2 ToM Agent’s Belief Structure of Multiple Agents
In our ToM framework, the focal agent maintains separate state beliefs ($s$) for itself and each other agent in the environment. In the case of a two-agent scenario, this yields $s=\{\underbrace{s^{f,\text{self}},s^{f,\text{world}}}_{\text{focal ($s^{f}$)}},\underbrace{s^{o,\text{self}},s^{o,\text{world}}}_{\text{other ($s^{o}$)}}\}$, where:
@@
This structure lets the focal agent keep its own perspective separate while modelling how others might see things differently. To elaborate, the belief components can be flexibly combined to capture different reasoning cases. For instance, the focal agent can pair $s^{o,\text{self}}$ with $s^{f,\text{world}}$ to predict what the other agent would observe given the focal agent’s own beliefs about the environment. The focal agent can also pair $s^{o,\text{self}}$ with $s^{o,\text{world}}$ to predict what the other agent thinks it will observe, based on the other agent’s own (possibly mistaken) view of the world. This kind of cross-perspective reasoning allows the focal agent to distinguish between (a) what it thinks the other agent will perceive and (b) what it thinks the other agent believes it may perceive. Because the focal agent’s own world beliefs ($s^{f,\text{world}}$) may differ from its beliefs about the other agent’s world beliefs ($s^{o,\text{world}}$), it can represent cases where knowledge is asymmetric – for example, when the focal agent knows something the other does not.
@@ 2.3 Recursive Planning with Theory of Mind
Step 1: Other Agent Policy Expansion. As aforementioned in Section 2.2, we begin with the focal agent’s beliefs, which comprises separate beliefs for itself and the other agent in the environment ($s=\{s^{f},s^{o}\}$). The focal agent first considers which policies the other agent is likely to select. This is visualised in Step 1 of Figure 1, where each policy node represents a specific action that the other agent might execute ($a_{0}$; purple square). The potential actions are evaluated based on the focal agent’s beliefs about the other agent’s beliefs ($s^{o}$; purple circle above purple squares). Essentially, the focal agent asks “Given what I believe about the other agent’s beliefs and goals, what would the other agent choose to do?”. The focal agent then computes how the other agent would update its beliefs if it were to execute that action ($s^{o}$; the purple circle below the purple square).
@@
Step 2: Focal Agent Policy Expansion. For each considered action of the other agent, the focal agent evaluates its own policy options. Crucially, before doing so, the focal agent updates its world beliefs, based on the anticipated consequences of the other agent’s actions, using likelihood message passing. Here, the focal agent uses its computation of how the other agent’s beliefs about the world ($s^{o,\text{world}}$; purple circle above the red circle) would change if the other agent were to execute an action. A likelihood message is then created, capturing the information gained from the other agent’s anticipated action in the form of the difference between the other agent’s updated beliefs and its prior beliefs. This mechanism allows the focal agent to incorporate information into its own beliefs about how the world states ($s^{f,\text{world}}$; red circle above the red squares) would change due to the other agent’s actions. Using these updated beliefs, the focal agent then evaluates its own policy options ($a_{0}$; red square) through standard EFE calculations. This creates branches in the tree structure for each possible joint policy combination between the focal and other agent. The focal agent then computes how its beliefs would be updated if it were to execute an action given the other agent’s action ($s^{f}$; red circle below the red square).
@@
Step 3: Focal Agent Observation Expansion. Then, given the joint policies, the focal agent considers what observations it is likely to receive and its resulting posterior beliefs.
@@
This process is illustrated in Step 3 of Figure 1, where the focal agent’s observation nodes ($o_{0}$; red diamonds) represent the various observations the focal agent expects to encounter given the focal agent’s beliefs considering the execution of both agents’ actions ($s^{f}$; red circle before red diamonds). This results in the computation of the focal agent’s posterior beliefs ($s^{f}$; red circle after red diamonds).
Step 4: Other Agent Observation Expansion. Here, the focal agent considers what observations the other agent is likely to receive ($o_{0}$; purple diamonds) given the joint policy and anticipated world state changes given its action. The observation probabilities are computed using the focal agent’s beliefs about the other agent’s self states ($s^{o,\text{self}}$; purple circle from an earlier expansion) and the focal agent’s own updated beliefs about the world state ($s^{f,\text{world}}$; red circle before the purple diamonds). The focal agent then updates its representation of the other agent’s posterior beliefs ($s^{o}$; purple circle after the purple diamond).
@@ 3.1 Collision Avoidance Task
Preferences encoded goal-seeking behaviour: high positive utility for reaching the target location and severe penalty for the null observation. Critically, no explicit collision avoidance preferences were included – coordination had to emerge through ToM reasoning rather than hard-coded behaviours. Planning horizons were set to 3 time steps for both ToM and non-ToM agents, sufficient to reach goals via alternative paths while requiring that much forward planning to identify coordination opportunities.
@@
In the ToM condition, the red (ToM) agent reasoned that the other agent will most likely move to the central location to take the shortest path to its goal, so it does not select moving to that location to avoid collision even though it was the most optimal policy for itself. Instead, the agent went around the centre, which was the next best alternative route towards its goal. The ToM agent selected a longer but collision-free route. For the detailed planning tree of the ToM vs non-ToM agents at time step 0, see Appendix 0.B.
@@ 3.2 Apple Foraging Task
The environment was partially observable such that agents could only assess apple availability at their current location, creating uncertainty about the broader resource distribution. Apples spawned probabilistically in orchard locations (25% per time step) and remained in that location until consumption. Action repertoires included movements (up, down, left, right actions), eating, and no operation. Preferences simply favoured reward acquisition without explicit cooperation incentives. Planning horizons were set to 3 time steps, sufficient for reaching the opposing end of the grid environment and consuming apples.
@@
For an illustration of the results from this task, see Figure 2(b). For the detailed planning tree of the ToM vs non-ToM agents at time step 0, see Appendix 0.C.
@@ Appendix 0.A Expected Free Energy under Sophisticated Inference with Theory of Mind
Appendix 0.B Planning Trees for Collision Avoidance Task
Figure 3 shows the difference in planning trees for both a non-ToM (c) and ToM agent (d) completing the collision avoidance task with another agent. It is clear that the ToM agent’s planning tree (d) entertains four routes the agent can take to avoid colliding with the other (non-ToM) agent and that the route illustrated in the right-most branch of the planning tree is selected to complete the task (a).
Figure 3: (a) Collision avoided by the red agent as the focal agent with ToM. (b) Reference grid layout for 3×3 simulation environment. Numbers 1-9 indicate cell indices used throughout the experiments to specify agent locations and movements. (c) Non-ToM planning tree: The red agent evaluates only its own policies over a 2-step horizon, selecting to go to location 5 (P==1.0), which eventually results in colliding with the other agent. (d) ToM planning tree: The red (ToM) agent recursively models the purple (non-ToM) agent’s policy space and beliefs, resulting four possible routes the red agent can take to avoid collision with the other agent which is predicted to go to its goal location via location 5.
Appendix 0.C Planning Trees for Apple Foraging Task
Figure 4 shows the difference in planning trees for both a non-ToM (c) and ToM agent (d) completing a apple foraging task with another agent. We start from the orchard environment state in (a), where both agents navigate to locate and consume apples, with initial shared knowledge of an apple at location 9. We visualise the planning tree for the red agent with (d) and without ToM (c).
In the non-ToM case, the focal agent (red) evaluates only its own policies over a 2-step horizon, selecting to go to location 9 (P==1.0) based on expected utility (G==10.00). However, when pursuing this policy it will end up with no apple, as the purple agent will get there first.
@@ Appendix 0.C Planning Trees for Apple Foraging Task
Figure 4: (a) Start situation with the red agent as the focal agent. (b) Reference grid layout for 3×3 simulation environment. Numbers 1-9 indicate cell indices used throughout the experiments to specify agent locations and movements. (c) Non-ToM planning tree: The red agent evaluates only its own policies over a 2-step horizon, selecting to go to location 9 (P==1.0) based on expected utility (G==10.00). (d) ToM planning tree: The red (ToM) agent recursively models the purple (non-ToM) agent’s policy space and beliefs, resulting in selection of going left with near full certainty (q_pi==0.95). This recursive reasoning leads to exploration of location 7, as the agent predicts the purple agent will pursue location 9, enabling coordinated foraging.